OpenStudy (anonymous):

The areas of two similar figures are 20 cm2 and 45 cm2. What is the scale factor of their corresponding sides? What is the ratio of their perimeters?

OpenStudy (mertsj):

\[\sqrt{\frac{20}{45}}=\sqrt{\frac{4}{9}}=\frac{2}{3}= scale factor of sides = ratio of perimeters\]

OpenStudy (anonymous):

Thanks!

OpenStudy (mertsj):

yw